Beavis Construction Company was the low bidder on a construction project to build an earthen dam for $1,890,000. The project was begun in 2017 and completed in 2018. Cost and other data are presented below:

2017 2018
Costs incurred during the year $ 423,000 $ 1,190,000
Estimated costs to complete 987,000 0
Billings during the year 445,000 1,445,000
Cash collections during the year 345,000 1,545,000

Assume that Beavis recognizes revenue on this contract over time according to percentage of completion.

Required:
Compute the amount of gross profit recognized during 2017 and 2018.


Beavis Construction Company was the low bidder on a construction project to build an earthen dam for $1,890,000. The project was begun in 2017 and completed in 2018. Cost and other data are presented below:

2017 2018
Costs incurred during the year $ 423,000 $ 1,190,000
Estimated costs to complete 987,000 0
Billings during the year 445,000 1,445,000
Cash collections during the year 345,000 1,545,000

Assume that Beavis recognizes revenue on this contract over time according to percentage of completion.

Required:
Compute the amount of gross profit recognized during 2017 and 2018.

Amount of gross profit recognized in 2017
Amount of gross profit recognized in 2018
Answer #1

Amount of gross profit recognized in 2017:

Revenue recognized = Total revenue - Cost till date

Revenue recognized = $1,890,000 - 1,410,000 (423,000+987,000) = $480,000

% of cost incurred during the year = $423,000 / 1,410,000 * 100 = 30%

Amount of gross profit recognized in 2017 = $480,000*30% = $144,000

Amount of gross profit recognized in 2018:

Revenue recognized = $1,890,000 - 1,613,000 (423,000+1,190,000) = $277,000

Amount of gross profit recognized in 2018 = $277,000 - 144,000 = $133,000